### v1

@@ 6 Evaluation
This section reports results from production and simulations on production data center traces.
6.1 End-to-end Simulation Results (LAS & LAVA)
We first compare performance of the different algorithms in simulation. These experiments are based on the C2 family. Our simulations are highly accurate.
Figure 6 shows simulation data from C2 production traces from 24 pools that cover a wide range of sizes, geographies, and usage patterns. The traces are from May-June 2024 and have a duration of over seven weeks. We measure the average percentage of empty hosts and report changes relative to our current production baseline. For instance, if LAVA achieves 21% and the baseline achieved 20%, we report a 1 percentage point (pp) improvement. We run all algorithms both with our actual lifetime model and a perfect (oracular) predictor as a comparison point. Similar to the LA paper (which increased packing density by 2% in production), a consistent 1 pp improvement represents a large gain in this scenario, and can be equivalent to 1% of a cluster’s capacity if used to power down or divest hosts, or offset holdbacks.
LAS outperforms LA-Binary on average, achieving an additional 1.1 pp increase in empty hosts. With more optimized misprediction mitigation, LAVA achieves best average improvement of 1.5 pp compared to LA-Binary.
We also observe that LAS consistently outperforms LA when both have oracular lifetimes: 9.5 pp vs. 7.5 pp on average, confirming our theoretical results that reprediction is fundamentally more powerful than one-shot prediction.
6.2 Production Results (LAS)
Prior to our fleet-wide roll-out, we ran several waves of production pilots on C2 and E2. For the first two waves, we conducted A/B experiments in three pools and measure the impact. We observe statistically significant empty host improvements ranging from 2% to 10% for the entire pool, shown in Table 1.
Table 1: Empty Host Improvements in Pilot Pools
@@ 6.2 Production Results (LAS)
For the third wave, we piloted the entire pool with LAS. We employ a Bayesian Structural Time Series method CausalImpact (Brodersen et al., 2015) to evaluate the pre-post causal effect brought by LAS (Figure 7). We observed a significant empty host improvement of 4.9% (Table 1). Moreover for the whole-pool pilot, we additionally measure the causal effect of LAS on stranded CPU and memory, where 1 pp of stranding reduction translates directly into 1% of capacity. LAS reduced CPU stranding by 3% (95% CI: [-5.5%, -0.49%]) and memory stranding by 2% (95% CI: [-3.8%, -0.23%]). We also conducted a whole-pool pilot for an E2 pool.
@@ 6.3 Lifetime-based Host Defragmentation (LARS)
Using production traces, we simulate defragmentation on two different one-month time intervals from a number of pools with customer workloads, including C2 and E2 VMs. In this experiment, we use oracle lifetimes. LARS does not affect which hosts are picked for migration at a given time, but orders the VMs across all in-migration hosts so that the VMs with the longest remaining lifetime are migrated first. The effect is that a larger number of VMs terminate before they would have been migrated, resulting in a reduction of around 4.5% of live migrations.
@@ 6.4 ML Model
Figure 7: Time series from our Wave 3 rollout, produced by the CausalImpact (Brodersen et al., 2015) package. Each sub-figure has three panels. The first panel shows the data and a counterfactual prediction (dotted line) for the post-launch period. The second panel shows the difference between observed data and counterfactual predictions, i.e., the point-wise causal effect, as estimated by the model. The third panel adds the point-wise contributions from the second panel, plotting the cumulative effect.
Our model achieves 70% recall at 99% precision, which is significantly higher than other model types we tried. We measured the latency of our model and find that most predictions complete in under 10 us (Figure 8). This latency improves upon Barbalho et al. (2023) by 780$\times$, and shows the benefit of compiling the model into the Borg Prime binary. Fast predictions make it practical to frequently re-run the model to correct mispredictions and to use it during maintenance and defragmentation.
@@ 6.5 Impact of Repredictions
To further understand the utility of repredictions leveraging the current VM uptime, we vary the amount of uptime used as input feature to the model and assess its impact on model accuracy. For each VM in the test trace, we divide its total lifetime into 20 quantiles. We then show the F1 score for predicting whether or not a VM lives beyond 168 hours (7 days) when the uptime is set to a particular quantile. Figure 9 illustrates the relationship between varying amount of uptime and model accuracy. We observe that without using reprediction (no uptime/0th quantile), the model only has a F1 score of 0.8. The F1 score quickly rises above 0.9 after the 8th quantile (i.e., once the VM’s uptime has reached 40% of its total lifetime). Observing how long a VM has lived, and repredicting its remaining lifetime thus leads to more accurate predictions. We find repredicting longer-lived VMs is critical to avoid spreading them across hosts and causing fragmentation.
@@ 7 Related Work
Prediction-based job scheduling has a long history. Some work uses RL Mao et al. (2019); Li et al. (2021); Wang et al. (2022) to learn decisions directly, such as for placing containers. Other work predicts properties of jobs and uses them to improve scheduling for scale-out Delimitrou & Kozyrakis (2014a), interference Delimitrou & Kozyrakis (2013), or goodput Jayaram Subramanya et al. (2023).
Job lifetimes are particularly useful for scheduling. Jockey Ferguson et al. (2012) and Aria Verma et al. (2011) predict runtimes in data processing frameworks and use them to improve scheduling of dataflow-based jobs. TetriSched Tumanov et al. (2016) uses runtime predictions in conjunction with an MILP solver to schedule jobs in an HPC-style setting. Wrangler uses them to identify stragglers Yadwadkar et al. (2014). 3Sigma Park et al. (2018) predicts various properties of Google job traces and uses them for scheduling. Note that most of this work uses runtime predictions in domain-specific scenarios and does not leverage general machine learning techniques to schedule.
In Cloud Computing, Microsoft’s Resource Central system Cortez et al. (2017) performs predictions on VMs and uses them for various tasks, including VM scheduling and admission control Sajal et al. (2023). Microsoft’s LA algorithm that performs on lifetime-based VM scheduling is the most similar to ours Barbalho et al. (2023); Buchbinder et al. (2022). Our work differs in that instead of performing one-shot predictions when the VM starts running, it repredicts VM lifetimes as needed, updating lifetimes based on new knowledge, delivering good improvements in empty hosts over LA. We show how to incorporate it in our existing scheduler by integrating the models into the binary. Our models run in-process at very low latency and our new scheduling algorithms tolerate mispredictions.

### v2

@@ 6 Evaluation
This section reports results from production and simulations on production data center traces. Our simulations are highly accurate (Appendix F). We report relative improvements and omit industry-competitive sensitive data.
6.1 End-to-end Simulation Results (NILAS & LAVA)
We first compare the performance of the different algorithms in simulation using C2 family production traces. Figure 6 shows simulation data from C2 traces from 24 pools that cover a wide range of sizes, geographies, and usage patterns. The traces are from May-June 2024 and have a duration of seven weeks. We measure the average percentage of empty hosts and report changes relative to the production baseline. For instance, if LAVA achieves 21% and the baseline achieved 20%, we report a 1 percentage point (pp) improvement. We run all algorithms both with our actual lifetime model and a perfect (oracular) predictor as a comparison point. Similar to the LA paper (which increased packing density by 2% in production), a consistent 1 pp improvement represents a large gain in this scenario, and can be equivalent to 1% of a cluster’s capacity if used to power down or divest hosts, or offset holdbacks.
NILAS outperforms LA-Binary on average, achieving an additional 1.1 pp increase in empty hosts. With more optimized misprediction mitigation, LAVA achieves the best average improvement of 1.5 pp compared to LA-Binary. We also observe that NILAS consistently outperforms LA when both have oracular lifetimes: 9.5 pp vs. 7.5 pp on average, confirming our theoretical results that reprediction is fundamentally more powerful than one-shot prediction.
As with many NP-hard problems, we observe occasional inversions when comparing techniques on large numbers of production pools with highly varying workloads. The different production pools we evaluated in Figure 6 vary significantly in size, utilization, VM workload distribution, and accuracy of the lifetime model. Isolating the specific factors driving these performance variations is challenging due to the complex interplay of these variables and direct correlations with individual factors were weak at best.
6.2 Production Results (NILAS)
Prior to fleet-wide rollout, we ran production pilots in waves on C2 and E2 traces. For the first two waves, we conducted A/B experiments in three pools and measured the impact. Table 2 shows our statistically significant empty host improvements, ranging from 2% to 10% for the entire pool.
Table 1: NILAS Empty Host Improvements in Pilot Pools
@@ 6.2 Production Results (NILAS)
For the third wave, we piloted NILAS on the entire pool. We employ a Bayesian Structural Time Series method CausalImpact (Brodersen et al., 2015) to evaluate the pre-post causal effect of NILAS (Figure 7). We observed a significant empty host improvement of 4.9% (Table 2). Moreover, for the whole-pool pilot, we additionally measure the causal effect of NILAS on stranded CPU and memory, where 1 pp of stranding reduction translates directly into 1% of capacity. NILAS reduced CPU stranding by 3% (95% CI: [-5.5%, -0.49%]) and memory stranding by 2% (95% CI: [-3.8%, -0.23%]). We also conducted a whole-pool pilot for an E2 pool. We note that C2 is slice-of-hardware while E2 is dynamically sized. The fact that the approach worked for both types of VM families is a non-obvious result, since dynamic resizing of VMs further complicates bin packing.
@@ 6.3 Lifetime-based Host Defragmentation (LARS)
Using production traces, we simulate defragmentation on two different one-month time intervals from a number of pools with customer workloads, including C2 and E2 VMs. In this experiment, we use oracle lifetimes. LARS does not affect which hosts are picked for migration at a given time, but orders the VMs on each in-migration host so that the VMs with the longest remaining lifetime are migrated first (Appendix H). The effect is that a larger number of VMs terminate before they would have been migrated, resulting in a reduction of around 4.5% of live migrations.
@@ 6.4 ML Model
Figure 7: Time series from our Wave 3 rollout, produced by CausalImpact (Brodersen et al., 2015). Each sub-figure has three panels. The first panel shows empty machines and a counterfactual prediction (dotted line) post-launch (labels omitted due to sensitivity). The second panel shows the difference between observed data and counterfactual predictions, i.e., the point-wise causal effect, as estimated by the model. The third panel adds the point-wise contributions from the second panel, plotting the cumulative effect.
Our lifetime model achieves 70% recall at 99% precision, which is significantly higher than other model types we tried (Appendix B). We measured the latency of our model and find that most predictions complete in under 10 us (Figure 9). This latency improves upon Barbalho et al. (2023) by 780$\times$, and shows the benefit of compiling the model into the Borg Prime binary. Fast predictions make it practical to frequently re-run the model to correct mispredictions and to use it during maintenance and defragmentation.
We analyze the impact of model features in Appendix A. We also analyze the sensitivity of NILAS and LAVA to prediction accuracy and show that our improvements apply across accuracies (Appendix  G.1). Finally, we performed a number of ablations to show that NILAS with oracle lifetimes and positioned high in the scoring function achieves close to the theoretical maximum of empty hosts (Appendix G.2). The gap between production performance and theoretical optimum is a combination of prediction accuracy and priority within the scoring function.
@@ 6.5 Impact of Repredictions
To further understand the utility of repredictions based on the current VM uptime, we vary the amount of uptime used as input feature to the model and assess its impact on model accuracy. For each VM in the test trace, we divide its total lifetime into 20 quantiles. We then show the F1 score for predicting whether or not a VM lives beyond 168 hours (7 days) when the uptime is set to a particular quantile. Figure 9 illustrates the relationship between varying amount of uptime and model accuracy. We observe that without using reprediction (no uptime/0th quantile), the model only has a F1 score of 0.8. The F1 score quickly rises above 0.9 after the 8th quantile (i.e., once the VM’s uptime has reached 40% of its total lifetime). Observing how long a VM has lived and repredicting its remaining lifetime thus leads to more accurate predictions. We find repredicting longer-lived VMs is critical to avoid spreading them across hosts and causing fragmentation.
One surprising result is that the the F1 score of quantiles 1-5 is, in fact, lower than the accuracy at the 0th quantile. This is because these uptimes are very close to 0 in absolute terms and thus difficult for the model to disambiguate, particularly since we are operating in the log domain. In contrast, the accuracy at 0% uptime is higher because it is always 0 and the model has seen many training samples with uptime=0. A potential optimization would be to only pass the uptime to the model if it reaches a particular threshold (e.g., 30 seconds), to avoid this initial dip in accuracy.
@@
6.6 Decrease in Model Accuracy Over Time
We find that the accuracy of our model remains stable for several months after deployment (Figure 10), but does require periodic retraining on the order of once a month to maintain its high accuracy. There are at least two reasons why accuracy drops over time. 1) New workloads arrive that the model was not trained on. 2) Workloads change their behavior. We looked at production data during this time period and observed gradual shifts in resource usage that is consistent with shifting workloads.
7 Production Experience
NILAS has been running in production since early 2024. We now share general insights from our experience of integrating ML into a mature production system. Our project started in 2020 and we quickly developed initial versions of our VM lifetime models and algorithms in simulation. Most time was spent on putting the approach in production.
Explainable models
One key requirement was that models had to be interpretable and explainable to be accepted in a production environment. We started with a lookup table approach where each entry contained a survival curve produced using Kaplan Meier Kaplan & Meier (1958). However, we found that the resulting models underperformed (Appendix B) and were not space-efficient. GBDTs showed the best combination of accuracy and explainability.
Model Deployment
We initially considered running the model within separate inference servers, but switched to an in-binary approach. Running our models at $<$10us latency became critical for our reprediction-based approach. We also found that depending on separate servers would have added a circular dependency from a reliability standpoint, since those servers would themselves run on top of Borg. For the same reason, we included the model in the binary as opposed to loading it from, e.g., a distributed file system.
@@ 7 Production Experience
Rollout
A major concern in our production environment was how to roll out the model. Rolling out the model independently of Borg risked the model becoming a single point of failure by bypassing existing verification approaches (e.g., gradual rollout and canary deployments). We found that treating the model like any other code change and rolling it out with new versions of the binary side-stepped this issue by subjecting models to the same careful production testing. We also relied extensively on simulations for backtesting and tuning design parameters, to avoid production changes.
Production Monitoring
During production rollout, we carefully monitored the behavior of NILAS, by adding additional telemetry and monitoring dashboards. This telemetry captures, e.g., the increase in CPU and memory caused by loading the model in production, the increase in time each scheduling pass takes, and how often the new cost component becomes the tie breaker. This instrumentation confirms the system is working as intended and debugs production issues. For example, in one of our initial deployments, we found a bug where the model was called incorrectly.
@@ 8 Related Work
Prediction-based job scheduling has a long history. Some work uses reinforcement learning Mao et al. (2019); Li et al. (2021); Wang et al. (2022), e.g., for placing containers. Other work predicts properties of jobs and uses them to improve scheduling for scale-out Delimitrou & Kozyrakis (2014a), interference Delimitrou & Kozyrakis (2013), or goodput Jayaram Subramanya et al. (2023).
Job lifetimes are particularly useful for scheduling. Jockey Ferguson et al. (2012) and Aria Verma et al. (2011) predict lifetimes (called runtimes) in data processing frameworks and use them to improve scheduling of dataflow-based jobs. TetriSched Tumanov et al. (2016) uses runtime predictions in conjunction with an MILP solver to schedule jobs in an HPC-style setting. Wrangler predicts stragglers Yadwadkar et al. (2014). 3Sigma Park et al. (2018) predicts various properties of Google job traces and uses them for scheduling. Note that most of this work uses runtime predictions in domain-specific scenarios and does not leverage general machine learning techniques to schedule.
In Cloud Computing, Microsoft’s Resource Central system Cortez et al. (2017) performs predictions on VMs and uses them for various tasks, including VM scheduling and admission control Sajal et al. (2023). Microsoft’s LA algorithm that performs lifetime-based VM scheduling is most similar to our work Barbalho et al. (2023); Buchbinder et al. (2021). Our work differs in that instead of performing one-shot predictions when the VM starts running, it repredicts VM lifetimes as needed, updating lifetimes based on new knowledge, delivering good improvements in empty hosts over LA, as shown in Section 6. We show how to incorporate lifetimes in our existing scheduler by integrating the models into the binary. Our models run in-process at very low latency and our new scheduling algorithms tolerate mispredictions. LAVA changes the objective function from LA and NILAS – instead of placing VMs with similar lifetimes on the same host, LAVA seeks to make more hosts empty by not extending the host lifetime, by placing short lived VMs on a host with one or more long lived VM(s).